### v2

# Assessing survival models by interval testing

Abstract
When considering many survival models, decisions become more challenging in health economic evaluation. In this paper, we present a set of methods to assist with selecting the most appropriate parametric survival models. The methods highlight areas of particularly poor fit. Furthermore, plots and overall p-values provide guidance on whether a parametric survival model should be rejected or not.
Keywords: Survival analysis, model selection, calibration, health economics, decision making
An R-package for these methods, including examples is available here.
1 Introduction
Right-censored survival data are frequently collected in clinical trials and often form the basis of a trial’s primary conclusions. To model the efficacy of the treatment within and beyond the duration of such trials, survival models can be fitted to such data. A choice between fitted survival models must then be made, and this is often the primary driving factor in the economic evaluation of a new medical product. During this process, a large number of models are usually considered. However, with a growing emphasis on flexible survival models [1] and a requirement for standard parametric models [2], the number of models considered can easily exceed 20, making decisions challenging. Our proposed methods can assist this process by indicating which survival models provide implausible estimates.
Recommendations on selecting the most appropriate survival models are provided in NICE TSD 14 [2]. These include visual curve inspection, AIC and BIC tests, clinical validity, and the use of external data. Except for AIC and BIC, these approaches are subjective and often time-consuming to implement, especially when considering many curves. It would be useful if certain models were quickly shown to be implausible. This would reduce the number of feasible models, making the decision process easier. One approach that may facilitate this is calibration, but this is rarely used in health technology assessment.
In survival analysis, calibration refers to the agreement between predicted probabilities and observed events within a specific time interval. The most common calibration methods are extensions of the Hosmer-Lemeshow goodness-of-fit test [3], [4]. These include the Grønnesby and Borgan test (specific to the Cox proportional hazard model) [5], the Nam and D’Agostino test (applicable to more general models) [6], as well as smoothed calibration curves and their associated metrics [7]. In all of these approaches, patients are assigned a risk score based on the fitted model and grouped (often into deciles) according to these scores. The groups are then evaluated to determine how well-calibrated a survival model is. Our approach is different from this because patients do not need to be assigned risk scores. This is especially useful for models where the risk scores are equal among all patients, such as in the standard parametric models recommended in NICE TSD 14 [2].
The paper is structured as follows: methods, application, simulation, and discussion. The methods section describes: a new calibration approach with intervals based on censoring, an extension to this where intervals can be specified, and methods for generating an overall p-value. The application section considers two examples. The first example evaluates a simple exponential model, providing an introduction to the developed methods. The second example is more practical and considers survival data which is challenging to model. In this example, seven parametric survival models are fitted, where the full range of developed methods are applied to evaluate which models are most appropriate. Subsequently, a simulation elucidates the validity of the various approaches, and type I error rates are presented for various data scenarios. Finally, a discussion is given.
2 Methods
2.1 Interval testing
We consider right-censored survival data, where for each subject we have (time, event) $\in(\mathbb{R}_{>0},\{1,0\})$. We split this data into two types:
  • The observed data (event=1) with times $t_{i,1}$ such that $0<t_{i,1}\leq t_{i+1,1}$ and $i=1,\ldots,E.$
  • The censored data (event=0) with times $t_{i,0}$ such that $0<t_{i,0}\leq t_{i+1,0}$ and $i=1,\ldots,C.$
Subsequently, we define the set of unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$. That is, the values $t_{j,0}^{\prime}$ are the unique values from the set of $t_{i,0}$ values, and thereby the $t_{j,0}^{\prime}$ values satisfy $0<t_{j,0}^{\prime}<t_{j+1,0}^{\prime}$. We now define the intervals of sequential censor times such that for j=1,2,,Jj=1,2,...,J
$I_{j}=\begin{cases}(0,\ t_{1,0}^{\prime}]\ \ \ \ \ \ \ \ \ for\ \ \ j=1\\ (t_{j-1,0}^{\prime},\ t_{j,0}^{\prime}]\ \ \ \ else\end{cases}$
We are interested in the hypothesis that the event generating process is described by the model $M(\boldsymbol{\beta})$, with probability density function $f_{M}$ and fixed parameters $\boldsymbol{\beta}$. We now make the assumption that the intervals themselves, $I_{j}$, are uninformative of the event generating process. It follows that under the null, the following process occurs within each interval
$events(I_{j})\sim binomial(N_{I_{j}},p_{I_{j}}),$ (1)
for cases where $N_{I_{j}}\geq 1$.
$events(I_{j})$ are the number of events observed within the interval $I_{j}$.
$p_{I_{j}}=$“the probability of a subject experiencing an event in interval $I_{j}$ given the subject makes it to interval $I_{j}$”, which is given by
$\begin{split}p_{I_{j}}&=\mathbb{P}(T\in I_{j}\ |\ T>min(I_{j})\ )\\ &=\mathbb{P}(T\in I_{j})\ /\ \mathbb{P}(T>min(I_{j}))\\ &=\int_{I_{j}}f_{M}(t)dt\Big/\int_{min(I_{j})}^{\infty}f_{M}(t)dt.\end{split}$ (2)
Finally, $N_{I_{j}}=$“the number of subjects eligible to experience an event at the start of interval $I_{j}$,” which we define for j=2,,Jj=2,...,J as
$\begin{split}&N_{I_{1}}=\ C+E\ -censors(I_{1}),\\ &N_{I_{j}}=\ C+E\ -\sum_{k=1}^{j-1}events(I_{k})\ -\sum_{k=1}^{j}censors(I_{k}),\end{split}$ (3)
where we define the number of censors observed within the interval $I_{k}$ as $censors(I_{k})$.
In summary, for intervals $I_{j}$ we have specified a binomial model with known $N_{I_{j}}$, $p_{I_{j}}$, and observed $events(I_{j})$. Each interval $I_{j}$, where $N_{I_{j}}\geq 1$, thereby has a midpoint p-value for observing $events(I_{j})$ in that interval.
2.2 Extension to specified intervals
The methodology above describes how events can be modelled within each interval. However, the intervals themselves are entirely determined by the censor times. It is often of interest to understand events in specified intervals (for example 10 evenly spaced intervals up to the last censor). Below, we extend the methodology to allow for such an analysis.
Define the intervals of interest for $k=1,\dots,K$ as
$V_{k}=\begin{cases}(S_{0},\ S_{1}]\ \ \ \ \ \ \ \ \ for\ \ \ k=1\\ (S_{k-1},\ S_{k}]\ \ \ \ \ \ else\end{cases}$
where $0<S_{0}<S_{1}<\dots<S_{K}$.
We wish to define a unique set of times $\tau_{i}$. We have unique censor times $t_{j,0}^{\prime}$ for $j=1,\ldots,J$ and unique $S_{k}$ values for $k=0,\dots,K$, taking the unique values of the union of these sets, we obtain the values $0<\tau_{i}<\tau_{i+1}$. From these, we construct the $I^{\prime}$ intervals for $i=1,2,...,L_{\tau}$, where $L_{\tau}$ denotes the total number of $\tau$ values.
$I^{\prime}_{i}=\begin{cases}(0,\tau_{1}]\ \ \ \ \ \ \ \ \ for\ \ \ i=1\\ (\tau_{i-1},\tau_{i}]\ \ \ \ \ \ else\end{cases}$
As in Equation 1, we now have a known distribution for each interval $I_{i}^{\prime}$. The observations in said intervals are described by
$events(I_{i}^{\prime})\sim binomial(N_{I_{i}^{\prime}},p_{I_{i}^{\prime}}),$
with $p_{I_{i}^{\prime}}$ and $N_{I_{i}^{\prime}}$ defined analogously to Equations 2 and 3, with all $I$ intervals in these formulae being replaced by $I^{\prime}$ intervals.
We now consider the $I_{i}^{\prime}$ intervals as subintervals that make up each $V_{k}$ interval. For each $V_{k}$, $k=1,\dots,K$, we define the collection of $L_{k}$ subintervals with $m=1,...,L_{k}$ as
$\{I_{m,V_{k}}^{\prime}\}=\{I_{i}^{\prime}\ \ \ such\ that\ \ I_{i}^{\prime}\subseteq V_{k}\}.$
It follows that for each interval $V_{k}$ its subintervals follow
[…]
Thereby $\sum^{L_{k}}_{m=1}events({I_{m,V_{k}}^{\prime}})$ is distributed as the realization of a sum of binomials with known parameters. That is, the number of observed events in each interval $V_{k}$ follows a known distribution. (In our code, we utilise that this is a special case of the Poisson binomial distribution.) We can thus obtain a midpoint p-value for each specified interval $V_{k}$.
2.3 Test statistics
The primary recommendation of this paper is the use of interval plots, such as Figure 1b. To support these plots, a single overall p-value can be a useful tool. Such a p-value can be obtained by condensing the plot’s constituent p-values into a single overall p-value. When deriving these test statistics, we assume that the p-value for each interval follows a continuous uniform[0,1] distribution under H0.
We derive two test statistics for obtaining an overall p-value. From Birnbaum (1954) [8], we know that selecting a best test statistic for H0 is not trivial, and a more modern discussion is given in [9]. Our first test statistic is continuous and allows for extreme p-values to have a high influence on the overall p-value. The second test statistic is discrete and aims to mitigate the influence of extreme p-values.
Transformed Fisher test (TFT) statistic
To motivate the test, we consider a p-value for a given interval. A high p-value indicates more events have occurred in the interval than would generally be expected under H0. A low p-value indicates fewer events than expected. It follows that both high and low p-values indicate poor model fit within a given interval. As such, we will construct a test statistic that can detect an uncommon amount of high and low p-values.
For each interval, we have a p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
$U_{j}=\begin{cases}2p_{j}\ \ \ \ \ \ \ \ \ \ \ for\ \ p_{j}\leq 0.5\\ 2(1-p_{j})\ \ \ \ for\ \ p_{j}>0.5\end{cases}$
It follows that $U_{j}\sim uniform[0,1]$. By the classical result from Fisher, we obtain the test statistic
$T_{cont}=-2\sum_{j=1}^{I}log(U_{j})\sim\chi^{2}_{{}_{2I}},$
where we reject for large $T_{cont}$ and $I$ is the total number of intervals. The intuition for this test is that unfavourable p-values (close to 0 or 1) are transformed such that they correspond to $U_{j}$ values close to 0. This makes the overall test able to detect unfavourable p-values.
Protection Against Very Small Intervals (PAVSI) test statistic
The transformed Fisher test is useful as it enables us to leverage the full range of reported p-values. However, in some situations, this may not be ideal. For example, a single p-value that is very close to 0 can heavily influence the test. We argue that a test that is not dominated by extreme p-values is a useful tool.
For each interval, we have a p-value, $p_{j}$. Under H0, we make the assumption these p-values are $uniform[0,1]$. The p-values are transformed to the form
$W_{j}=\begin{cases}1\ \ \ for\ \ p_{j}\geq 0.975\ \ or\ \ p_{j}\leq 0.025\\ 0\ \ \ else\end{cases}$
It follows that under H0 we have $W_{j}\sim bernoulli(0.05)$, and thereby
$T_{pavsi}=\sum_{j=1}^{I}W_{j}\sim binom(I,0.05),$
where we reject for large $T_{pavsi}$ values. It is noted that $T_{pavsi}$ is a discrete test statistic. Therefore, in Examples 1 and 2, the PAVSI is calculated using the midpoint p-value approach.
If the reader is interested in using PAVSI in there own work, there are some technical points to consider that are discussed in Appendix B.
3 Application
The methods above described two general approaches to interval testing:
  • Use the censor times to specify the intervals
  • Specify the intervals yourself (“extension to specified intervals”)
In example 1, we focus on the first approach only. In example 2, both approaches are described. For examples 1 and 2, we use real world overall survival data for melanoma from Freeman et al [10]. Example 1 uses the ‘Dabrafenib’ treatment arm of the BREAK-3 trial [11]. Example 2 uses the ‘Dabrafenib+Trametinib’ treatment arm of the COMBI-d trial [12]. All data and code is available on Github here.
3.1 Example 1: A fitted exponential model
In this example, we fit the simplest survival model (exponential) to our data. Following this, we want to determine how appropriate the exponential model is. First, we do a visual check of the curve fit, Figure 1a. Visually, the Kaplan-Meier curve and the exponential curve seem to be misaligned, suggesting poor fit. However, an interval test provides a more rigorous assessment of this, Figure 1b. For this example, we are only interested in an interval test where the intervals are selected using the censor times.
To perform the test, we first obtain the values $I_{j}$, $N_{I_{j}}$, and $events(I_{j})$ directly from the data. To complete equation 1, we must derive $p_{I_{j}}$ for the fitted exponential model. As the probability density function for the exponential distribution is given by $f_{M}(t)=\lambda e^{-\lambda t}$, using Equation 2, it follows that
$p_{I_{j}}=1-exp(\lambda[min(I_{j})-max(I_{j})]\ ).$
The value $\lambda$ is known (from fitting the exponential model to the data). We have therefore fully specified the binomial realisations for each interval $I_{j}$. This information for the first five intervals is presented in Table 1.
$I_{j}$ N.risk $p_{I_{j}}$ $events(I_{j})$ $\mathbb{E}$(events) p-mid
(0, 1] 187 0.026 1 4.8 0.028
(1, 1.6] 185 0.016 1 2.9 0.129
(1.6, 2.5] 183 0.024 1 4.4 0.038
(2.5, 3.4] 181 0.023 4 4.2 0.496
(3.4, 4.2] 176 0.021 5 3.6 0.775
Table 1: Example 1: Summary of interval data.
The p-value for each interval is discrete as it comes from a binomial model. It follows that each discrete p-value does not follow a continuous uniform[0,1] distribution under H0. This can be rectified by using randomised p-values [13]. However, in the interest of reproducibility and clarity, we will focus our discussion on the results of the midpoint p-values. (This is discussed in more detail in Appendix A.)
For each interval, we consider the midpoint p-value with respect to a 2-sided test. That is, if $p_{I_{j}}\leq 0.025$ or $p_{I_{j}}\geq 0.975$, we have reason to suspect the interval generating said value. We call this the individual test. Note that this is an informal tool, and we do accept or reject a model on the basis of a single individual test.
Of greater concern, are the intervals that fail the Bonferroni test [14]. The Bonferroni test is chosen to strongly control the familywise error rate at a level of 0.05. In particular, we reject if $p_{I_{j}}\leq 0.025/I$ or $p_{I_{j}}\geq 0.975/I$, where I=I=“the total number of intervals”. This means that under H0 (that the model holds for all intervals) a single Bonferroni rejection of an interval is only expected to occur for 1 in 20 times the data is generated under H0.
In Figure 1, we present the results for the Bonferroni and individual tests alongside the Kaplan-Meier (KM) curve, the fitted exponential model, and the numbers at risk. A single Bonferroni rejection of an interval is good reasoning for rejecting a model. In Figure 1, the very poor fit (Bonferroni rejection) is at the end of the curve. This is very unfavourable in health economic evaluation, where plausible extrapolations beyond the observed data are paramount to model selection. That is, the Bonferroni rejection at the end of the curve indicates that the extrapolation beyond the trial period may be poor.
An overall p-value for the fitted exponential model is also of interest. For the transformed Fisher test statistic, there are $I=42$ intervals, and our test statistic was computed as $2\sum_{j=1}^{I}log(U_{j})=81.15$, giving a p-value of 0.568. For the PAVSI test statistic, there are $I=42$ intervals. From Figure 1, we can read off that $T_{pavsi}=4$ (as there are 3 individual rejections and 1 Bonferroni rejection). This gives us an overall midpoint p-value of 0.107. (These test statistics should be used with caution as our simulation indicated that they are somewhat conservative in practice.)
As we can see, the two test statistics yield very different results. It is our recommendation that a model should not be accepted or rejected based entirely on the outcome of such tests. Instead, they are tools that provide an initial overview of model fit. The greatest detail and insight can be obtained by assessing the binomial interval plot alongside specific interval p-values and an overall p-value. For this particular example, we would recommend that the model is rejected. This is primarily due to the Bonferroni rejection at the end of the curve.
Figure 1: Example 1: Plot results. Part a: The KM curve for the data in purple, censors are given by dashed vertical black lines, and the fitted exponential curve is the overlaid curve in black. Part b: The binomial interval test for this data. Intervals are defined by the censors, where for each interval, red indicates a Bonferroni rejection, and grey indicates an individual flag. Part c: Numbers at risk for the data
3.2 Example 2: Selecting a parametric survival model for challenging data
In this example, the survival data has a challenging shape, making it difficult for models to provide good visual fit. We fit seven standard parametric survival models using flexsurv [15]. Our results and tables are then generated by the BITsurv package, which is available here. Our aim in Example 2 is to determine whether any of the seven survival models are appropriate, or whether survival models with additional flexibility are required. Both censor interval and specified interval approaches are used.
3.2.1 Intervals selected based on censors
First, we perform the interval test with intervals chosen using censors as in Example 1. The 7 parametric models were fitted, and the test statistics for each parametric model are presented in Table 2. The PAVSI and TFT overall p-values are well above the 0.05 threshold for all of the survival models. However, as will be discussed in our simulation section, these test statistics are often conservative for the censor interval approach.
For each model, there are between 1 and 4 individual flags. However, these flags are primarily descriptive. Of greater interest are the Bonferroni rejections, of which there are none. This seems good, and from Table 2 alone we might believe that our curves provide good fit. However, this is an unfortunate result of using intervals based on censors. Due to the large number of censors, the majority of intervals are very small, which makes it difficult to detect poor fit within these intervals. For example, there are many small intervals for the last 5 months of the curve. Instead, it would be more informative if we could consider a single interval for those final 5 months. To allow this, a specified interval approach is used.
Parametric survival model Bonferroni rejections Individual flags PAVSI TFT
Exponential 0 3 0.456 0.820
Gamma 0 2 0.686 0.859
Generalised gamma 0 1 0.876 0.939
Gompertz 0 4 0.257 0.853
Log-logistic 0 1 0.876 0.968
Log-normal 0 2 0.686 0.954
Weibull 0 3 0.456 0.841
Table 2: Example 2: Interval test results using censor defined intervals.
3.2.2 Ten evenly spaced intervals (a specified interval approach)
There are many ways these intervals could be chosen, and several approaches are briefly mentioned in the discussion. Here, we will outline one of the simpler approaches. The data is split into 10 evenly spaced intervals (owing to the classical convention of the Hosmer-Lemeshow test). These intervals start at time 0 and end at the largest censor time. That is, for k=1,,10k=1,...,10
$V_{k}=\frac{1}{10}\cdot t.max\cdot(k-1,\ k].$
Applying this approach to the data yields the results presented in Table 3. Of the seven fitted curves, we present the generalised gamma results in Figure 2.
The results of Table 3 tell a very different story to the censor interval approach. First, the overall TFT p-value is below 0.05 for all models except log-normal, indicating poor fit in 6 out of 7 models. Similarly, the overall PAVSI p-value mostly agrees with the TFT, rejecting 5 out of 7 models. However, the PAVSI should be interpreted with caution when using the 10-fixed interval approach as it can inflate the type I error rate, see discussion Appendix B. Finally, all of the parametric models have a Bonferroni rejection except log-normal. The log-normal model is described in greater detail in Appendix C, where the results for each of its 10 intervals are given in Table 6.
In conclusion, due to the high number of Bonferroni and TFT rejections, we suggest that the majority of the presented models are unsuitable. In terms of the interval tests, the log-logistic and log-normal were the best performers. However, even these models had their limitations, with the log-normal model being the worst of all 7 models for both AIC and BIC (not presented). For these reasons, it is recommended that exploring flexible models such as splines and piecewise would be appropriate for this data.
Parametric survival model Bonferroni rejections Individual flags PAVSI TFT
[…]
Table 3: Example 2: Interval test results using 10 evenly spaced intervals.
Figure 2: Example 2: Plot results for generalised gamma. Part a: The KM curve (purple), censors (vertical dashed lines), and the fitted generalised gamma curve (black). Part b: The binomial interval test results for 10 evenly spaced intervals. Part c: Numbers at risk.
4 Simulation
Due to the novelty of this approach, we will now pay particular attention to its properties under different simulated scenarios. Specifically, we focus on type I errors as this is the central concept about which the methodology is constructed.
4.1 Construction of simulation
We focus our attention on the exponential model as this allows us scope to explore other critical scenarios such as data maturity and patient numbers. Applications to additional scenarios (such as other parametric models) could be explored in future work.
Each simulation is described by the underlying event generation process: $T\sim exp(\lambda)$ and an underlying censor generating process for $C$ given by
$\begin{split}&C1\sim uniform[0,100],\\ &C2\sim uniform[18,22],\\ &C=min(C1,C2).\end{split}$
The observable trial data is constructed as $Time=min(C,T)$; and $Event=1$ if $T\leq C$, else $Event=0$. This construction can be visualised as a 22 Month trial with heavy censoring in the last 4 months of the trial and infrequent censoring throughout.
In total, 12 data scenarios will be explored. For the number of patients in the fitted treatment arm, we consider 4 possible sizes, $N:50,100,200$ and $500$. For data maturity, we consider three levels: $\lambda=1/10$ (mature), $\lambda=1/30$ (moderate), and $\lambda=1/70$ (immature), where at the end of the trial period (22 months), these levels of maturity imply underlying survival rates of $0.11,0.48,$ and $0.73$ respectively.
For all 12 scenarios, 10,000 simulations were run for each interval method (10 evenly spaced and censor) and each p-value approach (randomised and midpoint), totaling 480,000 simulated trial datasets. The results are presented here, and a rerun confirmed that any written interpretations remained consistent.
4.2 Simulation results: randomised p-values
A core assumption of our methodology is that the intervals are uninformative of the event-generating process. Randomised p-values allow us to test this core assumption via simulation, without the added concern of midpoint p-values influencing the results.
For all simulation scenarios, both the 10 fixed-interval and censor-interval approaches behaved well with respect to type I error rates. That is, these rates were close to 0.05 for both the Bonferroni and TFT tests. This indicates that the core assumption holds well in all of our simulated scenarios for both the 10 fixed-interval and censor-interval approaches. Tables for these results are provided in Appendix D.
4.3 Simulation results: midpoint p-values
Unlike randomised p-values, the midpoint p-values are not uniform[0, 1] under H0. Instead, this is an assumption we must make for practical purposes, see discussion in Appendix A. Below, we investigate the impact this assumption has on type I errors for our simulation.
The censor interval approach did not perform as well as the 10 fixed interval approach. Results for the censor interval approach are presented in Table 4. The type I error rates for the Bonferroni tests were slightly conservative, being between 0.01 and 0.04 in most cases. However, for the censor interval approach, the TFT was very conservative, having a type I error of 0.0001 in all the cases where $\lambda=1/30$ or $\lambda=1/70$. We argue that this occurs because a large proportion of each simulated dataset’s intervals had an expected number of events close to 0, and the majority of the observations in these intervals were 0. This results in an artificially high number of p-values close to 0.5 when using the midpoint p-value approach, influencing the TFT and causing the conservative results. (The 10 fixed interval approach does not suffer from this limitation.)
The 10 fixed interval approach performed preferably to the censor approach, with results generally being more stable and less conservative. Results for the 10 fixed interval approach are presented in Table 5. Data maturity had a moderate impact on the results. However, patient numbers had a more distinct impact, with lower patient numbers generally resulting in more conservative tests (lower type I errors). As patient numbers increased, the type I error rates appeared to approach the targeted level of 0.05. Across the scenarios, Bonferroni remained fairly stable (type I error rates generally between 0.02 and 0.05), whereas the TFT was slightly more variable (generally between 0.01 and 0.05).
4.4 Simulation conclusion
We explored the type I error rates of our methodology under a variety of scenarios for an underlying exponential survival model. As discussed in Appendix A, randomised p-values are not recommended for practical applications, but simulating this approach helped to validate the assumptions of our methodology. When midpoint p-values are used, both the 10 fixed interval and censor interval approaches were inclined to be conservative for all scenarios of our simulation. The 10 fixed interval approach was generally more stable, with type I error rates being closer to the targeted rate of 0.05. Because of this, we would, in general, recommend 10 fixed intervals as the preferred approach. However, the censor interval approach can still be a useful supporting tool.
Table 4: Simulation type I error results: Censor interval approach, midpoint p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0384 0.0253 0.0184 0.0001 0.0132 0.0001
100 0.0323 0.0091 0.0175 0.0001 0.0117 0.0001
200 0.0252 0.0009 0.0139 0.0001 0.0122 0.0001
500 0.0241 0.0001 0.0131 0.0001 0.0089 0.0001
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
Table 5: Simulation type I error results: 10 fixed interval approach, midpoint p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0325 0.0262 0.0219 0.0175 0.0245 0.0077
100 0.0390 0.0355 0.0336 0.0340 0.0221 0.0210
200 0.0436 0.0415 0.0421 0.0394 0.0336 0.0318
500 0.0491 0.0476 0.0440 0.0483 0.0410 0.0432
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
5 Discussion
Two approaches for interval testing were presented. In our first approach, intervals were selected by the data using the censor times. Unfortunately, this approach is generally unhelpful in describing regions with heavy censoring. This is because very small intervals are unlikely to be rejected due to low power. Additionally, with this approach, our simulations indicated that the test statistics are conservative in standard cases. However, we argue that the censor interval plot is still a useful tool. Specifically, by visually inspecting the censor interval plot and test results, one is more supported when informally deducing complex properties of the data.
In the second approach, we extended the methodology so that intervals can be specified. Specifically, we focused on the 10 evenly spaced interval approach. The 10 intervals were generally easier to interpret with plots compared to the censor interval approach. Additionally, the approach is informative for areas with heavy censoring, such as the tail of the KM curve, which is often the area of greatest importance for long-term extrapolation. Furthermore, our simulations indicated that this method generally performed better than the censor interval approach, especially if the number of patients exceeded 100. (That is, the type I errors were closer to the targeted 0.05 level.) For these reasons, we would generally recommend the 10 evenly spaced interval approach to the censor interval approach, especially for cases where the approaches are being used to formally accept or reject fitted survival models. To facilitate the application of these approaches, a package has been developed.
Regarding further research, there are several avenues for further exploration. Although we focused on 10 evenly spaced intervals, the methodology is flexible to accommodate many methods of interval selection. One possible extension would be automatic interval selection based on equal statistical power. Another extension could involve restricting chosen intervals to have boundaries that are censor times. Another avenue of exploration is to apply our methodology to risk score models. That is, more complex survival models that allow unequal risk scores for each patient (our examples inherently assumed equal risk scores). This approach would require some modifications to the methodology, specifically using a Poisson binomial distribution rather than the binomial distribution (Equation 1).
Our approach provides several advantages over standard model selection techniques such as AIC/BIC. First, the TFT and Bonferroni provide objective tests, which can be used to reject a fitted survival model. Moreover, the methods indicate areas of particularly poor fit (for example, any one of 10 evenly spaced intervals up to the end of the KM curve). Such information is invaluable to analysts, helping them to decide which survival models are most suitable.
Appendix A Midpoint p-values discussion
When deriving test statistics and results, we will often make the assumption that midpoint p-values are uniform[0,1] under H0. If the validity of this approach is in question, then randomized p-values can be used, which are known to be uniform[0,1] under H0 [13]. For this reason, the randomized p-values are superior from this purely statistical standpoint. However, we argue below that the midpoint p-values are more practical and that, for our applications, the uniform[0,1] assumption is often reasonable.
In the case of reducing the number of intervals to 10, as is described in example 2, the generating process is the sum of binomials. If the number of patients is reasonable, one would expect these intervals to contain a reasonable number of feasible discrete outcomes. This means the uniform[0,1] assumption is more valid compared to when there are only a small number of possible outcomes.
We also acknowledge that in practice, randomized p-values are more difficult to deal with. When an analysis is rerun, the p-values are rerandomised. This would then involve rerunning the analysis to determine the level of variation in the randomised p-values and their resulting test statistic. This is helpful. However, this approach leads you down the path of using the expected value of these reruns as your final quoted value. That is, the midpoint p-value, which is the expected value of the randomised p-value.
Appendix B PAVSI - practical discussion
In this paper, we focused primarily on the TFT test statistic as it utilises the full range of the data. That is, for PAVSI, a p-value of say 0.026 would contribute equivalently to a p-value of 0.5, which is arguably a limitation of this approach compared to the TFT. For this reason, in general, we recommend the TFT over the PAVSI.
Additionally, there is a caveat to using the PAVSI for the 10 evenly spaced interval approach when using midpoint p-values. Specifically, even if the p-values are uniform[0,1] under H0, when we use the midpoint p-value to obtain the PAVSI of the 10 evenly spaced intervals, it can be shown analytically that the type I error is 0.08614 rather than the targeted 0.05 level. Because of this limitation, the PAVSI results are not described in the simulation section. Regardless, interested readers can find these simulation results on Github. Additionally, PAVSI is included as an option in the available codes.
In conclusion, for the censor interval approach, the PAVSI behaves well and can be well used. For the 10 evenly spaced interval approach, the PAVSI should be used with caution as the type I error rate can be inflated. In real-world analyses, extreme observations in an interval (that is likely very small) can render the TFT useless on account of a single extreme p-value, which results in the TFT rejecting almost every model. For such cases, PAVSI is the recommended alternative. For standard cases, the TFT is generally preferred.
Appendix C Additional details for Example 2: Log normal model with 10 evenly spaced intervals
In Example 2, we looked at seven fitted survival models using a 10 evenly spaced interval approach. Of those seven models, only the log-normal model was not rejected by the Bonferroni test. Because of this, we explore the log-normal model in further detail. Specifically, we now look at the results in each of the 10 intervals for the log-normal model, Table 6.
From Table 6, we see that the only interval resulting in an individual flag for the log-normal model was between times 30.35 and 34.69 with a p-value of 0.0061, which is above the 0.0025 threshold of a Bonferroni rejection. Furthermore, we note that the 4th interval had a p-value of 0.9727. If this value was only 0.0023 higher, then the log-normal would have received a second individual flag, and as a result of this the PAVSI overall p-value would have yielded a rejection (noting caveat Appendix B). Furthermore, the log-normal performed the worst of all 7 models for both AIC and BIC (not presented). For these reasons, as well as the overall TFT p-value of 0.0513 (very close to 0.05) we can be justifiably suspicious of the log-normal model. However, we have no strong evidence to outright reject it.
Table 6: Example 2: Interval test details for the log-normal model with 10 evenly spaced intervals.
$V_{j}$ $\mathbb{E}$(Events) Obs Events p-mid Individual test Bonferroni test
(0, 4.34] 18.67 11 0.0301 Accept Accept
(4.34, 8.67] 22.55 23 0.5514 Accept Accept
(8.67, 13.01] 17.83 26 0.9690 Accept Accept
(13.01, 17.34] 13.59 21 0.9727 Accept Accept
(17.34, 21.68] 10.70 11 0.5549 Accept Accept
(21.68, 26.01] 8.99 7 0.2550 Accept Accept
(26.01, 30.35] 7.69 9 0.6975 Accept Accept
(30.35, 34.69] 6.51 1 0.0061 Flag Accept
(34.69, 39.02] 4.95 5 0.5364 Accept Accept
(39.02, 43.36] 1.56 3 0.8598 Accept Accept
Appendix D Simulation results when using randomised p-values
As mentioned previously, the type I error rates for both Tables 7 and 8 were close to 0.05 for both the Bonferroni and TFT tests. However, it is noted that for the censor interval approach with $\lambda=1/10$ and $N=50$ the TFT and Bonferroni rates were slightly higher being close to 0.07, Table 7. This may indicate that for small N with a high rate of events our assumption (that the intervals are uninformative of the event‐generating process) may not hold as well compared to other cases.
Table 7: Simulation type I error results: Censor interval approach, randomised p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0692 0.0713 0.0557 0.0584 0.0500 0.0524
100 0.0535 0.0635 0.0513 0.0541 0.0465 0.0562
200 0.0547 0.0601 0.0521 0.0540 0.0521 0.0526
500 0.0531 0.0557 0.0493 0.0511 0.0494 0.0499
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.
Table 8: Simulation type I error results: 10 fixed interval approach, randomised p-values
$\boldsymbol{\lambda=1/10}$ $\boldsymbol{\lambda=1/30}$ $\boldsymbol{\lambda=1/70}$
Number of patients Bonferroni TFT Bonferroni TFT Bonferroni TFT
50 0.0510 0.0550 0.0528 0.0561 0.0527 0.0549
100 0.0537 0.0518 0.0502 0.0504 0.0523 0.0554
200 0.0494 0.0543 0.0512 0.0556 0.0470 0.0539
500 0.0512 0.0492 0.0474 0.0487 0.0456 0.0532
For each of the 12 scenarios (data maturity and patient numbers), 10,000 datasets were simulated.